A logging barge on the Mississippi River travels 24 miles upstream in 3 hours. The return trip takes the barge only 2 hours. Find the rate of the barge in still water.
Answer by mananth(16946)

let speed of barge in still water be x mph
and current speed = y mph
upstream distance = 24
time = 3 hours
upstream speed = 24/3 = 8 mph
downstream distance = 24 miles
time = 2 hours
speed = 24/2=12 mph

x+y =12
x-y =8
add the equations
2x = 20
x= 10
speed of bargein still water = 10 mph